a cashier has a total of 125 bills made up of fives and tens. The total value of the money is $815. How many ten-dollar bills does the cashier have?
Answer by nerdybill(7384)

a cashier has a total of 125 bills made up of fives and tens. The total value of the money is $815. How many ten-dollar bills does the cashier have?
.
Let x = number of ten-dollar bills
then from "a cashier has a total of 125 bills" we know:
125-x = number of five-dollar bills
.
Since the total is 815:
10x + 5(125-x) = 815
10x + 625 - 5x = 815
5x + 625 = 815
5x = 190
x = 38